OpenStudy (anonymous):

What is the probability that out of 125 babies born, at least 70 will be girls? Assume that boys and girls are equally probable, and round your answer to the nearest tenth of a percent. A.99.0% B.70.4% C.10.5% D.51.2%

OpenStudy (anonymous):

10 20 30 40 50 60 70 / 80 90 100 110 120 ...125 70 / 50... Which I pick... B.

OpenStudy (anonymous):

By the way, I made up that way to get the answer, not a real technique.

OpenStudy (amistre64):

im thinking it sounds like a binomial probability setup

OpenStudy (anonymous):

that what these are

OpenStudy (amistre64):

125 C 70 times success^(70) times failure^(125-70) 125 C 70 (1/2)^(70) (1/2)^(125-70) soo:\[\sum_{k=70}^{125}\binom{125}{k}.5^{k}.5^{125-k}\]

OpenStudy (amistre64):

or we can try a normal approximation :)

OpenStudy (amistre64):

mean = np = 125(.5) = 62.5 so we would expect 62.5 of the population to be girls sd = sqrt(npq) = sqrt(62.5*.5) = sqrt(31.25)

OpenStudy (amistre64):

do you have to do this all by hand or do you have a ti83?

OpenStudy (amistre64):

\[z=\frac{x-mean}{sd}\] \[z=\frac{70-62.5}{\sqrt{31.25}}=1.34\] looking on a ztable; that corresponds to about 100% - 91%

OpenStudy (amistre64):

so yeah, it aint gonna be B, but it does look like it will be close to 10%